Complete the table below for contribution margin per unit, total contribution margin, and contribution margin ratio:


Compute the missing information, starting with scenario A, then for scenarios B and C. (Enter the contribution margin ratio to nearest percent, X%.) 
Answer #1
A B C
Number of units 1290 Units 13860 Units 2620 Units
Sale price per unit 1400 4300 5330
Variable cost per unit 700 2150 3731
Calculate
Contribution margin per unit 700 4300-2150 = 2150 5330-3731 = 1599
Total Contribution margin 700*1290 = 903000 2150*13860 = 29799000 1599*2620 = 4189380
Contribution margin ratio 50% 50% 30%